# Negative non-Gaussianity as a salvager for PBHs with PTAs in bounce

### v1

@@ Appendix A Comoving curvature perturbation mode solutions
In this section, we underline the explicit curvature perturbation mode solutions for each phase in our EFT of bounce set-up. These solutions will later help us to construct the tree-level scalar power spectrum and are needed further to evaluate the various momentum and temporal integrals for the quantum loop corrections. The discussions related to one-loop quantum corrections are part of the next section. The upcoming results are basically solutions of the Fourier space MS equation, see Eqn. (9), for any wavenumber $k$. Our complete setup consists of $5$ phases, starting with the contraction, followed by a bounce into an expanding phase, which joins with the SRI as inflation begins, a sharp transition into the USR, and a sharp transition into the final SRII phase where inflation ends. For both contracting and bouncing phases, we do not restrict to any particular matter or ekpyrotic kind but generalize the study to suit either.
@@
    where we notice the appearance of two new sets of Bogoliubov coefficients namely, $\alpha_{2},\;\beta_{2}$. These coefficients can be calculated keeping in mind the above two formulas in Eqn. (82-83) and applying the continuity and differentiability conditions for the modes at the instant $\tau=\tau_{s}$ of the transition. Solving these boundary conditions would yield the following solutions for the coefficients:
@@
    where, via the similar continuity and differentiabilty condition applied to the solution in Eqn. (83-86), provides with the following new set of coefficients:
@@
Appendix B One-loop corrections using In-In formalism
@@
The above relations combine, after using the mode solutions for each phase in our set-up, to give the corresponding total correction at the one-loop and thus we can write this sum from Eqs. (91,92,93) as:
@@
Appendix C Regularized-Renormalized-Resummed formulas
@@
After the renormalization, the late-time secular growth of the remaining IR divergences, coming in the form of Logarithms of different wavenumbers from each phase, can be softened further via the DRG resummation procedure (see III.2.3). The final result uses the terms in Eqn. (97-101) after evaluating them at the pivot scale ()(*)( ∗ ) and they package the effect of IR divergences in the form of quantum corrections which reads:

### v3

@@ Appendix B Comoving curvature perturbation mode solutions
In this Appendix, we underline the explicit curvature perturbation mode solutions for each phase in our EFT of bounce set-up. These solutions will later help us construct the tree-level scalar power spectrum and are needed further to evaluate the various momentum and temporal integrals for the quantum loop corrections. The discussions related to one-loop quantum corrections are part of the next section. The upcoming results are basically solutions of the Fourier space MS equation, see Eqn. (9), for any wavenumber $k$. Our complete setup consists of $5$ phases, starting with the contraction, followed by a bounce into an expanding phase, which joins the SRI as inflation begins, a sharp transition into the USR, and a sharp transition into the final SRII phase where inflation ends. For both contracting and bouncing phases, we do not restrict to any particular matter or ekpyrotic kind but generalize the study to suit either.
@@
    where we notice the appearance of two new sets of Bogoliubov coefficients namely, $\alpha_{2},\;\beta_{2}$. These coefficients can be calculated keeping in mind the above two formulas in Eqn. (92-93) and applying the continuity and differentiability conditions for the modes at the instant $\tau=\tau_{s}$ of the transition. Solving these boundary conditions would yield the following solutions for the coefficients:
@@
    where, via the similar continuity and differentiabilty condition applied to the solution in Eqn. (93-96), provides with the following new set of coefficients:
@@
Appendix C One-loop corrections using In-In formalism
@@
The above relations combine, after using the mode solutions for each phase in our set-up, to give the corresponding total correction at the one-loop and thus we can write this sum from Eqs. (101,102,103) as:
@@
Appendix D Regularized-Renormalized-Resummed formulas
@@
After the renormalization, the late-time secular growth of the remaining IR divergences, coming in the form of Logarithms of different wavenumbers from each phase, can be softened further via the DRG resummation procedure (see III.2.3). The final result uses the terms in Eqn. (107-111) after evaluating them at the pivot scale ()(*)( ∗ ) and they package the effect of IR divergences in the form of quantum corrections which reads: